# homogeneous_khovanskii.bib

### v1

@@ 3.1. MUVAK bases
Let $f_{1},\dots,f_{k}\in A$ be $\Delta$/̄homogeneous generators and let $f\in A$ be a $\Delta$/̄homogeneous element. We say that $f$ is faithfully representable in $f_{1},\dots,f_{k}$ with respect to $v_{1},\dots,v_{m}$, if there is a polynomial $h\in K[X_{1},\dots,X_{k}]$ such that $h(f_{1},\dots,f_{k})=f$ and for all $t\in\operatorname{Supp}(h)$ and for all $i\in\{1,\dots,m\}$, we have $v_{i}(t(f_{1},\dots,f_{k}))\succeq v_{i}(f)$.
@@
Let $\mathcal{B}\subseteq A$ be a set of $\Delta$/̄homogeneous generators of $A$ as a $K$/̄algebra. We say that $\mathcal{B}$ is a $\Delta$/̄homogeneous MUVAK basis(1)(1)(1)Multiple Valuations Analogue to Khovanskii bases of $A$ with respect to $v_{1},\dots,v_{m}$, if for every $\Delta$/̄homogeneous $f\in A$ there are $f_{1},\dots,f_{k}\in\mathcal{B}$, such that $f$ is faithfully representable in $f_{1},\dots,f_{k}$ with respect to $v_{1},\dots,v_{m}$.
While we take full responsibility for inventing the name, $\Delta$/̄homogeneous MUVAK bases already appeared in the literature in the construction of certain Cox rings. We give more details on this application in Section 5.
We relate MUVAK bases to Khovanskii bases. For this, as in the previous section, we let $f_{1},\dots,f_{k}\in A$ be $\Delta$/̄homogeneous generators of $A$ as a $K$/̄algebra and consider the morphism
@@
and this gives $\Gamma$/̄gradings via $\deg_{i}(X_{j})\coloneqq v_{i}(f_{j})$ for all $i\in\{1,\dots,m\}$. The morphisms $\alpha$ and $\beta_{i}$ are $\Delta$/̄graded by construction and $\beta_{i}$ is a $\Gamma$/̄graded morphism via $\deg_{i}$.
Notation 3.3.
We write $\operatorname{in}_{i}(h)\in K[X_{1},\dots,X_{k}]$ for the sum of the terms of minimal $\deg_{i}$/̄degree of a polynomial $h\in K[X_{1},\dots,X_{k}]$ (with respect to $\succeq$).
@@
The set $\{f_{1},\dots,f_{k}\}$ is a $\Delta$/̄homogeneous MUVAK basis of $A$ with respect to $v_{1},\dots,v_{m}$ if and only if for every $\Delta$/̄homogeneous $f\in A$ there is a $\Delta$/̄homogeneous $h\in K[X_{1},\dots,X_{k}]$ with $\alpha(h)=f$ and $\beta_{i}(\operatorname{in}_{i}(h))=\operatorname{in}_{v_{i}}^{\Delta}(f)$ for all $i\in\{1,\dots,m\}$.
@@
Let $f\in A$ be $\Delta$/̄homogeneous. If $\{f_{1},\dots,f_{k}\}$ is a $\Delta$/̄homogeneous MUVAK basis with respect to $v_{1},\dots,v_{m}$, then $f$ must be faithfully representable in $f_{1},\dots,f_{k}$, so there is $h\in K[X_{1},\dots,X_{k}]$ with $\alpha(h)=f$ and $\deg_{i}(\operatorname{in}_{i}(h))=v_{i}(f)$ for all $i\in\{1,\dots,m\}$. The maps $\beta_{i}$ are $\deg_{i}$/̄graded, so we must have $\beta_{i}(\operatorname{in}_{i}(h))=\operatorname{in}_{v_{i}}^{\Delta}(f)$ for degree reasons.
Conversely, assume that there is a $\Delta$/̄homogeneous $h\in K[X_{1},\dots,X_{k}]$ with $\alpha(h)=f$ and $\beta_{i}(\operatorname{in}_{i}(h))=\operatorname{in}_{v_{i}}^{\Delta}(f)$ for all $i\in\{1,\dots,m\}$. Then $\deg_{i}(\operatorname{in}_{i}(h))=\deg_{v_{i}}(\operatorname{in}_{v_{i}}(f))=v_{i}(f)$ as required. ∎
@@
Let $v:B\setminus\{0\}\to\Gamma$ be a valuation over $K$ and let $f_{1},\dots,f_{k}\in A$ be $\Delta$/̄homogeneous generators. If $\{f_{1},\dots,f_{k}\}$ is a $\Delta$/̄homogeneous MUVAK basis of $A$ with respect to $v$, then $\{f_{1},\dots,f_{k}\}$ is a $\Delta$/̄homogeneous Khovanskii basis of $A$ with respect to $v$.
Conversely, if $\{f_{1},\dots,f_{k}\}$ is a $\Delta$/̄homogeneous Khovanskii basis of $A$ with respect to $v$ and Algorithm 1 terminates for $A$, then $\{f_{1},\dots,f_{k}\}$ is a $\Delta$/̄homogeneous MUVAK basis of $A$ with respect to $v$.
@@
$\beta:K[X_{1},\dots,X_{k}]\to\operatorname{in}_{v}^{\Delta}(A),\leavevmode\nobreak\ X_{i}\mapsto\operatorname{in}_{v}^{\Delta}(f_{i})$
@@
Assume that $\{f_{1},\dots,f_{k}\}$ is a $\Delta$/̄homogeneous MUVAK basis and let $f\in A$ be $\Delta$/̄homogeneous. By Lemma 3.4, there is a $\Delta$/̄homogeneous polynomial $h\in K[X_{1},\dots,X_{k}]$ with $\alpha(h)=f$ and $\beta(\operatorname{in}_{\Gamma}(h))=\operatorname{in}_{v}^{\Delta}(f)$. Hence $\beta$ is surjective and $\{f_{1},\dots,f_{k}\}$ is a $\Delta$/̄homogeneous Khovanskii basis.
Assume now that $\{f_{1},\dots,f_{k}\}$ is a $\Delta$/̄homogeneous Khovanskii basis or, equivalently, that $\beta$ is surjective. Let $f\in A$ be $\Delta$/̄homogeneous. We run Algorithm 1 for $f$ together with $\{f_{1},\dots,f_{k}\}$. By assumption, the algorithm terminates, and by surjectivity of $\beta$, the algorithm must give remainder $r=0$ and an element $h\in K[X_{1},\dots,X_{k}]$ with $f=h(f_{1},\dots,f_{k})$ and $\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h))=v(f)$. Hence $f$ is faithfully representable in $f_{1},\dots,f_{k}$ with respect to $v$ via $h$. ∎
@@
It appears that being a MUVAK basis is a stronger property than being a Khovanskii basis. It is not clear what an analogue for the filtration given by one valuation in the case of multiple valuations should be. Finding such an analogue appears to be crucial for fully integrating MUVAK bases into the theory of SAGBI and Khovanskii bases.
@@ 3.2. Negative homogenization
To give an algorithm for the computation of a $\Delta$/̄homogeneous MUVAK basis, we require the notion of ‘negative homogenization’.
Let $R\coloneqq K[X_{1},\dots,X_{k}]$ be a polynomial ring graded by a matrix $W\in\mathbb{Z}^{m\times k}$ of rank $m\geq 1$, that is, $\deg_{W}(X_{j})\coloneqq(W_{1j},\dots,W_{mj})\in\mathbb{Z}^{m}$, $1\leq j\leq k$. We introduce additional variables $t_{1},\dots,t_{m}$ and endow the ring $R^{-}\coloneqq K[X_{1},\dots,X_{k},t_{1},\dots,t_{m}]$ with the grading induced by the matrix $W^{-}\coloneqq(W\mid-I_{m})$, where $I_{m}$ denotes the identity matrix of rank $m$. Analogous to the homogenization with respect to variables of degree 1, see for example [KR05, Definition 4.3.1], we define the homogenization of a polynomial $f\in R$ with respect to the variables $t_{1},\dots,t_{m}$ of degree $-1$.
@@
    For the zero polynomial, we set $00^{\mathrm{hom}}=$ 0.
@@
Write $f=f_{1}+\cdots+f_{s}$ with $f_{i}$ the terms of $f$. Let $\mu_{+}\coloneqq\operatorname{topdeg}_{-W}(f)$ (see [KR05, Definition 4.3.1]) and $\mu_{-}\coloneqq\operatorname{mindeg}_{W}(f)$. We observe that $-\mu_{+}=\mu_{-}$. We have
@@
Lemma 3.8 enables us to use well/known ‘calculation rules’ for homogenized polynomials and ideals as in [KR05, Proposition 4.3.2] and [KR05, Proposition 4.3.5], although these rules are only proven for homogenizations with respect to variables of positive degree in the given references.
@@ 3.3. Computing homogeneous MUVAK bases
Yamagishi [Yam18] gives an algorithm for the computation of a $\Delta$/̄homogeneous MUVAK basis in the special situation discussed in section 5. In the following, we present an algorithm that appears to perform better than Yamagishi’s algorithm. However, just as in [Yam18], the algorithm is restricted to the case $\Gamma=\mathbb{Z}$. From now on, we hence assume that $\Gamma=\mathbb{Z}$ with the natural ordering $\geq$.
Let $f_{1},\dots,f_{k}\in A$ be $\Delta$/̄homogeneous generators of $A$ as a $K$/̄algebra and consider the morphisms $\alpha$ and $\beta_{i}$ for $i\in\{1,\dots,m\}$ as above. Let $R\coloneqq K[X_{1},\dots,X_{k}]$. We collect the gradings $\deg_{i}$ on $R$ in a matrix $W\in\mathbb{Z}^{m\times k}$ via $W_{ij}\coloneqq\deg_{i}(X_{j})$. In the following, the homogenization operator $(\cdot)^{\mathrm{hom}}$ always means the negative homogenization with respect to the grading given by $W$ (mapping from $R$ to $R[t_{1},\dots,t_{m}]$). We extend the gradings $\deg_{i}$ to $R[t_{1},\dots,t_{m}]$ by putting $\deg_{i}(t_{j})=-\delta_{ij}$ for $1\leq i,j\leq m$, where $\delta_{ij}$ denotes the Kronecker delta. Let $I\coloneqq\ker(\alpha)$.
We have the following characterization of faithful representability.
@@
Let $f\in A$ be $\Delta$/̄homogeneous, let $h\in\alpha^{-1}(f)$ be any preimage and put
@@
Let $f$ be faithfully representable and let $h^{\prime}\in\alpha^{-1}(f)$ such that for all $t\in\operatorname{Supp}(h^{\prime})$ and for all $i\in\{1,\dots,m\}$, we have $v_{i}(t(f_{1},\dots,f_{k}))\geq v_{i}(f)$. If $v_{i}(f)=\deg_{i}(h^{\mathrm{hom}})$, then $\vartheta=1$ and $h^{\mathrm{hom}}\in I^{\mathrm{hom}}+\langle 1\rangle$ is trivially fulfilled. So, assume that $\deg_{i}(h^{\mathrm{hom}})<v_{i}(f)$ for some $i\in\{1,\dots,m\}$. As $\alpha(h)=\alpha(h^{\prime})$, we have $h-h^{\prime}\in I$, so $(h-h^{\prime})^{\mathrm{hom}}\in I^{\mathrm{hom}}$. By assumption, we have $\deg_{i}(h^{\prime\mathrm{hom}})=v_{i}(f)$, hence $(h-h^{\prime})^{\mathrm{hom}}=h^{\mathrm{hom}}-h^{\prime\mathrm{hom}}\vartheta$. Hence
@@
Conversely, let $h^{\mathrm{hom}}\in I^{\mathrm{hom}}+\langle\vartheta\rangle$. Then there are $h^{\prime}\in R[t_{1},\dots,t_{m}]$ and $h^{\prime\prime}\in I^{\mathrm{hom}}$ with $h^{\mathrm{hom}}=h^{\prime\prime}+h^{\prime}\vartheta$. We may assume that $h^{\prime}$ is $\deg_{i}$/̄homogeneous for all $i\in\{1,\dots,m\}$ with $\deg_{i}(h^{\prime})=\deg_{i}(h^{\mathrm{hom}})-\deg_{i}(\vartheta)=v_{i}(f)$. We hence have $\deg_{i}(t)\geq v_{i}(f)$ for all $t\in\operatorname{Supp}(h^{\prime\mathrm{deh}})$. Further $h-h^{\prime\mathrm{deh}}\in I$ by [KR05, Proposition 4.3.5 b)], so $f$ is faithfully representable via $h^{\prime\mathrm{deh}}$. ∎
Notation 3.11.
For $h\in K[X_{1},\dots,X_{k}]$ and $i\in\{1,\dots,m\}$, we write
@@
Let $f\in A$ be $\Delta$/̄homogeneous. If for all $h\in\alpha^{-1}(f)$ and for all $i\in\{1,\dots,m\}$ we have $\operatorname{mindeg}_{i}(h)=v_{i}(f)$ or $\operatorname{in}_{i}(h)^{\mathrm{hom}}\in I^{\mathrm{hom}}+\langle t_{i}\rangle$, then $f$ is faithfully representable in $f_{1},\dots,f_{k}$ with respect to $v_{1},\dots,v_{m}$.

### v2

@@ 3.1. MUVAK bases
Let $f_{1},\dots,f_{k}\in A$ be $\Delta$-homogeneous generators and let $f\in A$ be a $\Delta$-homogeneous element. We say that $f$ is faithfully representable in $f_{1},\dots,f_{k}$ with respect to $v_{1},\dots,v_{m}$, if there is a polynomial $h\in K[X_{1},\dots,X_{k}]$ such that $h(f_{1},\dots,f_{k})=f$ and for all $t\in\operatorname{Supp}(h)$ and for all $i\in\{1,\dots,m\}$, we have $v_{i}(t(f_{1},\dots,f_{k}))\succeq v_{i}(f)$.
@@
Let $\mathcal{G}\subseteq A$ be a set of $\Delta$-homogeneous generators of $A$ as a $K$-algebra. We say that $\mathcal{G}$ is a $\Delta$-homogeneous MUVAK basis(1)(1)(1)Multiple Valuations Analogue to Khovanskii bases of $A$ with respect to $v_{1},\dots,v_{m}$, if for every $\Delta$-homogeneous $f\in A$ there are $f_{1},\dots,f_{k}\in\mathcal{G}$, such that $f$ is faithfully representable in $f_{1},\dots,f_{k}$ with respect to $v_{1},\dots,v_{m}$.
While we take full responsibility for inventing the name, $\Delta$-homogeneous MUVAK bases already appeared in the literature in the construction of certain Cox rings. We give more details on this application in Section 5.
Example 3.3.
We again illustrate the results of this section with a running example, which is continued in 3.15, 3.18 and 3.22. Let $K$ be a field, $B=K[x,y]$ and $v_{x}$ and $v_{y}$ be the order of divisibility by $x$ and $y$, respectively, with the natural ordering on $\mathbb{Z}$. Let
$f_{1}\coloneqq x^{2}+xy,f_{2}\coloneqq x^{2},f_{3}\coloneqq y^{2}+xy,f_{4}\coloneqq y^{2}\in B$
and set $A=K[f_{1},f_{2},f_{3},f_{4}]$, so $A$ is the subalgebra generated by all monomials of even degree. For simplicity, we do not consider a grading in this example and put $\Delta=\{0\}$ (although everything works with the standard grading by $\Delta=\mathbb{Z}$). We have the following valuations and initial forms:
$\begin{array}[]{c|c|c}&v_{x}&v_{y}\\ \hline\cr f_{1}&1&0\\ f_{2}&2&0\\ f_{3}&0&1\\ f_{4}&0&2\end{array}\quad\quad\quad\begin{array}[]{c|c|c}&\operatorname{in}_{v_{x}}&\operatorname{in}_{v_{y}}\\ \hline\cr f_{1}&xy&x^{2}\\ f_{2}&x^{2}&x^{2}\\ f_{3}&y^{2}&xy\\ f_{4}&y^{2}&y^{2}\end{array}$
We claim that $\{f_{1},\dots,f_{4}\}$ is not a MUVAK basis of $A$ with respect to $v_{x},v_{y}$. Consider $f=f_{1}-f_{2}=xy\in A$. For degree reasons, any polynomial $h\in K[X_{1},\dots,X_{4}]$ with $h(f_{1},f_{2},f_{3},f_{4})=xy$ must be of the form $h=a_{1}X_{1}+a_{2}X_{2}+a_{3}X_{3}+a_{4}X_{4}$ with $a_{i}\in K$. Furthermore, we must have $a_{1}\neq 0$ or $a_{3}\neq 0$ because only the polynomials $f_{1}$ and $f_{3}$ involve the monomial $xy$. If $a_{1}X_{1}\in\operatorname{Supp}(h)$, then $v_{y}(a_{1}f_{1})<v_{y}(xy)=1$, and if $a_{3}X_{3}\in\operatorname{Supp}(h)$, then $v_{x}(a_{3}f_{3})<v_{x}(xy)=1$. Hence $f$ is not faithfully representable in $f_{1},\dots,f_{4}$ with respect to $v_{x}$ and $v_{y}$.
Before we present an algorithm for the computation of MUVAK bases, we relate MUVAK bases to Khovanskii bases, see Theorem 3.6 for the precise statement. For this, as in the previous section, we let $f_{1},\dots,f_{k}\in A$ be $\Delta$-homogeneous generators of $A$ as a $K$-algebra and consider the morphism
@@
and this gives $\Gamma$-gradings via $\deg_{i}(X_{j})\coloneqq v_{i}(f_{j})$ for all $i\in\{1,\dots,m\}$. The morphisms $\alpha$ and $\beta_{i}$ are $\Delta$-graded by construction and $\beta_{i}$ is a $\Gamma$-graded morphism via $\deg_{i}$.
Notation 3.4.
We write $\operatorname{in}_{i}(h)\in K[X_{1},\dots,X_{k}]$ for the sum of the terms of minimal $\deg_{i}$-degree of a polynomial $h\in K[X_{1},\dots,X_{k}]$ (with respect to $\succeq$).
We have the following equivalent characterization of a MUVAK basis using the maps $\alpha$ and $\beta_{i}$.
@@
The set $\{f_{1},\dots,f_{k}\}$ is a $\Delta$-homogeneous MUVAK basis of $A$ with respect to $v_{1},\dots,v_{m}$ if and only if for every $\Delta$-homogeneous $f\in A$ there is a $\Delta$-homogeneous $h\in K[X_{1},\dots,X_{k}]$ with $\alpha(h)=f$ and $\beta_{i}(\operatorname{in}_{i}(h))=\operatorname{in}_{v_{i}}^{\Delta}(f)$ for all $i\in\{1,\dots,m\}$.
@@
Let $f\in A$ be $\Delta$-homogeneous. If $\{f_{1},\dots,f_{k}\}$ is a $\Delta$-homogeneous MUVAK basis with respect to $v_{1},\dots,v_{m}$, then $f$ must be faithfully representable in $f_{1},\dots,f_{k}$, so there is $h\in K[X_{1},\dots,X_{k}]$ with $\alpha(h)=f$ and $\deg_{i}(\operatorname{in}_{i}(h))=v_{i}(f)$ for all $i\in\{1,\dots,m\}$. The maps $\beta_{i}$ are $\deg_{i}$-graded, so we must have $\beta_{i}(\operatorname{in}_{i}(h))=\operatorname{in}_{v_{i}}^{\Delta}(f)$ for degree reasons.
Conversely, assume that there is a $\Delta$-homogeneous $h\in K[X_{1},\dots,X_{k}]$ with $\alpha(h)=f$ and $\beta_{i}(\operatorname{in}_{i}(h))=\operatorname{in}_{v_{i}}^{\Delta}(f)$ for all $i\in\{1,\dots,m\}$. Then $\deg_{i}(\operatorname{in}_{i}(h))=\deg_{v_{i}}(\operatorname{in}_{v_{i}}(f))=v_{i}(f)$ as required. ∎
@@
Let $v:B\setminus\{0\}\to\Gamma$ be a valuation over $K$ and let $f_{1},\dots,f_{k}\in A$ be $\Delta$-homogeneous generators. If $\{f_{1},\dots,f_{k}\}$ is a $\Delta$-homogeneous MUVAK basis of $A$ with respect to $v$, then $\{f_{1},\dots,f_{k}\}$ is a $\Delta$-homogeneous Khovanskii basis of $A$ with respect to $v$.
Conversely, if $\{f_{1},\dots,f_{k}\}$ is a $\Delta$-homogeneous Khovanskii basis of $A$ with respect to $v$ and Algorithm 1 terminates for $A$, then $\{f_{1},\dots,f_{k}\}$ is a $\Delta$-homogeneous MUVAK basis of $A$ with respect to $v$.
@@
$\beta:K[X_{1},\dots,X_{k}]\to\operatorname{in}_{v}^{\Delta}(A),\penalty 10000\ X_{i}\mapsto\operatorname{in}_{v}^{\Delta}(f_{i})$
@@
Assume that $\{f_{1},\dots,f_{k}\}$ is a $\Delta$-homogeneous MUVAK basis and let $f\in A$ be $\Delta$-homogeneous. By Lemma 3.5, there is a $\Delta$-homogeneous polynomial $h\in K[X_{1},\dots,X_{k}]$ with $\alpha(h)=f$ and $\beta(\operatorname{in}_{\Gamma}(h))=\operatorname{in}_{v}^{\Delta}(f)$. Hence $\beta$ is surjective and $\{f_{1},\dots,f_{k}\}$ is a $\Delta$-homogeneous Khovanskii basis.
Assume now that $\{f_{1},\dots,f_{k}\}$ is a $\Delta$-homogeneous Khovanskii basis or, equivalently, that $\beta$ is surjective. Let $f\in A$ be $\Delta$-homogeneous. We run Algorithm 1 for $f$ together with $\{f_{1},\dots,f_{k}\}$. By assumption, the algorithm terminates, and by surjectivity of $\beta$, the algorithm must give remainder $r=0$ and an element $h\in K[X_{1},\dots,X_{k}]$ with $f=h(f_{1},\dots,f_{k})$ and $\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h))=v(f)$. Hence $f$ is faithfully representable in $f_{1},\dots,f_{k}$ with respect to $v$ via $h$. ∎
@@
It appears that being a MUVAK basis with respect to a single valuation is a stronger property than being a Khovanskii basis. It is not clear what an analogue for the filtration given by one valuation in the case of multiple valuations should be. Finding such an analogue appears to be crucial for fully integrating MUVAK bases into the theory of SAGBI and Khovanskii bases.
Remark 3.8.
If $\{f_{1},\dots,f_{k}\}$ is a $\Delta$-homogeneous MUVAK basis with respect to $v_{1},\dots,v_{m}$, then $\{f_{1},\dots,f_{k}\}$ is a $\Delta$-homogeneous MUVAK basis with respect to every single valuation $v_{i}$ as well. Hence $\{f_{1},\dots,f_{k}\}$ is a $\Delta$-homogeneous Khovanskii basis with respect to every valuation $v_{i}$ by Theorem 3.6.
The converse statement is in general wrong: The set $\{f_{1},\dots,f_{4}\}$ in Example 3.3 is a Khovanskii basis for $A$ with respect to $v_{x}$ as well as with respect to $v_{y}$, but it is not a MUVAK basis with respect to $v_{x}$ and $v_{y}$.
@@ 3.2. Negative homogenization
To give an algorithm for the computation of a $\Delta$-homogeneous MUVAK basis, we require the notion of ‘negative homogenization’.
Let $R\coloneqq K[X_{1},\dots,X_{k}]$ be a polynomial ring graded by a matrix $W\in\mathbb{Z}^{m\times k}$ of rank $m\geq 1$, that is, $\deg_{W}(X_{j})\coloneqq(W_{1j},\dots,W_{mj})\in\mathbb{Z}^{m}$, $1\leq j\leq k$. We introduce additional variables $t_{1},\dots,t_{m}$ and endow the ring $R^{-}\coloneqq K[X_{1},\dots,X_{k},t_{1},\dots,t_{m}]$ with the grading induced by the matrix $W^{-}\coloneqq(W\mid-I_{m})$, where $I_{m}$ denotes the identity matrix of rank $m$. Analogous to the homogenization with respect to variables of degree 1, see for example [KR05, Definition 4.3.1], we define the homogenization of a polynomial $f\in R$ with respect to the variables $t_{1},\dots,t_{m}$ of degree 1-1.
@@
    For the zero polynomial, we set $0^{\mathrm{hom}}=0$.
@@
Write $f=f_{1}+\cdots+f_{s}$ with $f_{i}$ the terms of $f$. Let $\mu_{+}\coloneqq\operatorname{topdeg}_{-W}(f)$ (see [KR05, Definition 4.3.1]) and $\mu_{-}\coloneqq\operatorname{mindeg}_{W}(f)$. We observe that $\mu_{+}=\mu_{-}$. We have
@@
Lemma 3.10 enables us to use well-known ‘calculation rules’ for homogenized polynomials and ideals as in [KR05, Proposition 4.3.2] and [KR05, Proposition 4.3.5], although these rules are only proven for homogenizations with respect to variables of positive degree in the given references.
@@ 3.3. Computing homogeneous MUVAK bases
Yamagishi [Yam18] gives an algorithm for the computation of a $\Delta$-homogeneous MUVAK basis in the special situation discussed in section 5. In the following, we present an algorithm that appears to perform better than Yamagishi’s algorithm. However, just as in [Yam18], the algorithm is restricted to the case $\Gamma=\mathbb{Z}$. From now on, we hence assume that $\Gamma=\mathbb{Z}$ with the natural ordering $\geq$.
Let $f_{1},\dots,f_{k}\in A$ be $\Delta$-homogeneous generators of $A$ as a $K$-algebra and consider the morphisms $\alpha$ and $\beta_{i}$ for $i\in\{1,\dots,m\}$ as above. Let $R\coloneqq K[X_{1},\dots,X_{k}]$. We collect the gradings $\deg_{i}$ on $R$ in a matrix $W\in\mathbb{Z}^{m\times k}$ via $W_{ij}\coloneqq\deg_{i}(X_{j})$. In the following, the homogenization operator $(\cdot)^{\mathrm{hom}}$ always means the negative homogenization with respect to the grading given by $W$ (mapping from $R$ to $R[t_{1},\dots,t_{m}]$). We extend the gradings $\deg_{i}$ to $R[t_{1},\dots,t_{m}]$ by putting $\deg_{i}(t_{j})=-\delta_{ij}$ for $1\leq i,j\leq m$, where $\delta_{ij}$ denotes the Kronecker delta.
Notation 3.12.
For $h\in K[X_{1},\dots,X_{k}]$ and $i\in\{1,\dots,m\}$, we write
We observe:
@@
Let $f\in A$ be $\Delta$-homogeneous and let $h\in\alpha^{-1}(f)$. Then
$\deg_{i}(h^{\mathrm{hom}})=\operatorname{mindeg}_{i}(h)\leq v_{i}(f)$
for all $i\in\{1,\dots,m\}$.
Proof.
The first equality holds because $(\cdot)^{\mathrm{hom}}$ denotes a negative homogenization. The inequality follows from the definition of the gradings $\deg_{i}$. ∎
Let $I\coloneqq\ker(\alpha)$. We have the following characterization of faithful representability.
Lemma 3.14.
Let $f\in A$ be $\Delta$-homogeneous, let $h\in\alpha^{-1}(f)$ be any preimage and put
@@
Let $f$ be faithfully representable and let $h^{\prime}\in\alpha^{-1}(f)$ such that for all $t\in\operatorname{Supp}(h^{\prime})$ and for all $i\in\{1,\dots,m\}$, we have $v_{i}(t(f_{1},\dots,f_{k}))\geq v_{i}(f)$. As $\alpha(h)=\alpha(h^{\prime})$, we have $h-h^{\prime}\in I$, so $(h-h^{\prime})^{\mathrm{hom}}\in I^{\mathrm{hom}}$. By assumption, we have $\deg_{i}(h^{\prime\mathrm{hom}})=v_{i}(f)$ and by Lemma 3.13, we have $\operatorname{mindeg}_{i}(h)\leq\operatorname{mindeg}_{i}(h^{\prime})$ for all $i\in\{1,\dots,m\}$. Hence $(h-h^{\prime})^{\mathrm{hom}}=h^{\mathrm{hom}}-h^{\prime\mathrm{hom}}\vartheta$ and therefore
@@
Conversely, let $h^{\mathrm{hom}}\in I^{\mathrm{hom}}+\langle\vartheta\rangle$. Then there are $h^{\prime}\in R[t_{1},\dots,t_{m}]$ and $h^{\prime\prime}\in I^{\mathrm{hom}}$ with $h^{\mathrm{hom}}=h^{\prime\prime}+h^{\prime}\vartheta$. We may assume that $h^{\prime}$ is $\deg_{i}$-homogeneous for all $i\in\{1,\dots,m\}$ with $\deg_{i}(h^{\prime})=\deg_{i}(h^{\mathrm{hom}})-\deg_{i}(\vartheta)=v_{i}(f)$. We hence have $\deg_{i}(t)\geq v_{i}(f)$ for all $t\in\operatorname{Supp}(h^{\prime\mathrm{deh}})$. Further $h-h^{\prime\mathrm{deh}}\in I$ by [KR05, Proposition 4.3.5 b)], so $f$ is faithfully representable via $h^{\prime\mathrm{deh}}$. ∎
Example 3.15.
This example is part of the series also containing 3.3, 3.18 and 3.22. Let again $f=xy$. The homogenization matrix is given by
$W=\begin{pmatrix}1&2&0&0\\ 0&0&1&2\end{pmatrix}$.
Computing the relations of $f_{1},\dots,f_{4}$ gives
$I=\langle X_{1}-X_{2}-X_{3}+X_{4},X_{3}^{2}-X_{2}X_{4}-2X_{3}X_{4}+X_{4}^{2}\rangle$
with homogenization
$I^{\mathrm{hom}}=\langle X_{1}t_{1}-X_{2}t_{1}^{2}-X_{3}t_{2}+X_{4}t_{2}^{2},X_{3}^{2}-X_{2}X_{4}t_{1}^{2}-2X_{3}X_{4}t_{2}+X_{4}^{2}t_{2}^{2}\rangle.$
For the polynomial $h=X_{1}-X_{2}$ with $h(f_{1},\dots,f_{4})=f$, we have $h^{\mathrm{hom}}=X_{1}-X_{2}t_{1}$, $\deg_{1}(h^{\mathrm{hom}})=1$ and $\deg_{2}(h^{\mathrm{hom}})=0$. Hence $\vartheta=t_{2}$ and we see that $h^{\mathrm{hom}}\notin I^{\mathrm{hom}}+\langle t_{2}\rangle$. So, Lemma 3.13 confirms that $f$ is not faithfully representable in $f_{1},\dots,f_{4}$.
@@
Let $f\in A$ be $\Delta$-homogeneous. If for all $h\in\alpha^{-1}(f)$ and for all $i\in\{1,\dots,m\}$ we have $\operatorname{mindeg}_{i}(h)=v_{i}(f)$ or $\operatorname{in}_{i}(h)^{\mathrm{hom}}\in I^{\mathrm{hom}}+\langle t_{i}\rangle$, then $f$ is faithfully representable in $f_{1},\dots,f_{k}$ with respect to $v_{1},\dots,v_{m}$.